Solve the following frame using Slope Deflection Method,

30 kN/m A B1 Ic D 12 kN/m 3m 12 kN/m + 2m 4m am
Answer #1

Sol B 30kn/m 3m 12 kN/m E AF 42 m & um- D2m Assuming Flexural zooting Rigidity (EI) = constant for all membe Fix end moments2) Now slope deflection Requation is given as. for Span АВ, MAQ M Ag + ЧЕТ ЗА 262 98 6CI A р МА — ма+ 4Єгоо ТЕто - 6 Ето e apfor span CD Now directly applying valus McD = esh O 2 0+ UEIOC 2 EI OD 2 Moc = o it Y EIQp + 2 EI OG o eso 2 for span EB- o eFrom can 1 USIQA - 2A OB 7 2 20A DB DA= - OB 2 foom er 6, 4 EA OD =-265 Oc Oc 12 from est Eac o > 2 EI OB 8 Soy 2 EI 20€ + 0B(6 Mg, C -3ч. 38 3 см 0.ҫет 90+3.5 ЄI. = -32- + from cu 15 & 16, we get SIOе = 33 EI Oct — 33 Зо Now applying these vabes inMcg= Чо+ Еz 9 + CI Op население - 3 ч. 383 Зох2. Mce = 2ч 383 km) 2 €т 9. + Мер = 2610 + CI Op с + вт хө Стхэс долоон = 2єID